# BATIS: Bootstrapping, Autonomous Testing, and Initialization System for Quantum Dot Devices

### v1

@@ Abstract
Semiconductor quantum dot (QD) devices have become central to advancements in spin-based quantum computing. As the complexity of QD devices grows, manual tuning becomes increasingly infeasible, necessitating robust and scalable autotuning solutions. Tuning large arrays of QD qubits depends on efficient choices of automated protocols. Here, we introduce a bootstrapping, autonomous testing, and initialization system (BATIS), an automated framework designed to streamline QD device testing and initialization. BATIS navigates high-dimensional gate voltage spaces, automating essential steps such as leakage testing and gate characterization. The current channel formation protocol follows a novel and scalable approach that requires a single measurement regardless of the number of channels. Demonstrated at on a quad-QD Si/SixGe1-x device, BATIS eliminates the need for deep cryogenic environments during initial device diagnostics, significantly enhancing scalability and reducing setup times. By requiring minimal prior knowledge of the device architecture, BATIS represents a platform-agnostic solution, adaptable to various QD systems, which bridges a critical gap in QD autotuning.
@@ I Introduction
Quantum dot (QD) qubit devices are commonly formed by electrostatically manipulating the potential energy landscape of a two-dimensional (2D) electron or hole gas (EG or HG). This procedure produces interleaved regions of accumulated and depleted charge density that serve as QD qubits, charge sensors, and charge carrier reservoirs [1, 2, 3, 4, 5]. In recent years, the number of QDs in such devices has risen rapidly [6, 7, 8]. In the near term, human intervention will no longer be possible to tune up many dots simultaneously. The field of autotuning has arisen to address this problem.
@@
Modern semiconductor QD qubit devices have two critical features that enhance their performance and simultaneously make autotuning challenging and essential. First, the active region of the device—the QDs, charge sensors, and reservoirs—contains no doping. All charge carriers—electrons or holes—are induced by gate voltages. No current can flow until gate voltages are applied and carriers are accumulated, yet those voltages are unknown a priori. Moreover, the exact voltage configuration changes not only between cooldowns but even between successive initialization steps. Second, QD device chips have multiple current channels that often serve different purposes. Some can host an array of QD qubits, while others provide a home for charge sensors. These channels are not separate until pinch-off voltages, i.e., the gate voltage threshold at which carrier accumulation is eliminated, are determined. Moreover, since the accumulation and depletion of carriers must occur in alternating regions separated by only tens of nanometers to enable the formation of QDs separated by tunnel barriers and proximal charge sensors, these channels must interact.
@@
Here, we describe a bootstrapping, autonomous testing, and initialization system (BATIS) that can automatically test the device and set the gate voltages to ensure that the QD chip is properly accumulated and all required pinch-off voltages are determined. BATIS takes as input the device gate layout and a high-level understanding of the desired operation, including critically where current channels should be continuous and which gates should divide one channel from another. The outputs of BATIS are the gate voltage parameters that make the QD chip functional, with all current channels well delineated and all pinch-off voltages known. Importantly, in this work, the sample is kept at around –about 10 times warmer than other typical devices [9, 10]. Such temperatures are highly desirable for what is known as high-temperature operation [11]. However, higher temperature causes an increased difficulty in thresholding pinchoffs as the coulomb blockade and all measurement features derived therein are less well defined and easier to confuse with noise [12]. BATIS is capable of tuning at these high temperatures, which is important both for scalability and automated testing as, for example, is possible using cryogenic wafer probers [13, 14].
@@
Previous work on autotuning has shown that a bootstrap process is sufficient for further tuning of QD devices. For example, if current channels are known in modulation-doped QD devices, the formation of double QDs with few-electron occupation can be automated [15]. In such devices, autonomous bootstrapping has been demonstrated [16, 17]. If a charge sensor is available, undoped SiMOS devices have been autotuned to achieve a single-electron QD [18]. For nanowires, which by design have a single, known current channel, automation into fully formed QDs has been demonstrated, provided device-specific details are placed into a configuration file [19].
@@
For accumulation-mode, undoped QD devices, previous work has enabled progress in tuning device topology [20, 21] and charge-state calibration [16, 22]. Methods for automated detection of unintentional, spurious QDs have also been proposed [23]. More recently, a framework to achieve orthogonal control of the electrostatic potential landscape in a large 2D array has been put forward [24]. However, autonomous bootstrapping for accumulation mode QD devices has never been attempted, and most tuning algorithms proposed to date assume that steps such as device characterization and the definition of the QD(s) channel(s) have already been completed [25].
The approach for designing BATIS differs from others, typically exercised by the community. All bootstrapping algorithms proposed to date are highly device-specific, with the algorithm’s overall efficiency measured as progress-per-second or some other metric against time used as a figure of merit. However, time-based metrics for device-specific algorithms provide little basis for valid comparison against alternative methods. Such metrics quickly become instrument-bound. Taking large numbers of data points per measurement, each with integration times in the or , quickly dwarfs the simplest point-wise threshold analysis performed on modern multi-core CPUs clocked in the . Users can easily speed up their benchmarks for single-use cases by upgrading to faster instrumentation.
Moreover, the resulting algorithms provide single-use, device-specific applications by assuming substantial knowledge about the tuned state of QD devices. Thus, even if such specialized algorithms perform well on the development device, they require significant redesigning to be deployable on other QD devices.
In contrast, when creating BATIS, we focused on developing a platform-agnostic tool that can be applied to a variety of QD devices. While the procedural flow of BATIS is predetermined, the executable flow is automatically adjusted based on the general device architecture and connectivity to the measurement instruments, both provided within the configuration files. In other words, rather than being programmed to follow a fixed sequence of measurements, BATIS is designed to choose which steps to take and change the expected outcome of each step – depending on the device used. Further adjustments are automatically made as the tuning progresses and the algorithm learns the device.
The device characterization and tuning processes are a repetitive yet essential part of initializing QD-based qubit chips. The majority of autotuning efforts have focused on the latter phases of automatic device tuning, including coarse tuning, charge-state tuning, virtualization, and fine-tuning, which is sufficient for quantum dots in a one-dimensional channel or when the initial step is performed by hand [25]. In modern chips, the scale of the tune-up will soon preclude hand-tuning as fully 2D structures become ubiquitous. BATIS addresses the challenges that are about to become limiting in this widespread class of quantum semiconductor chips.
@@ II BATIS Design
Figure 1(a) shows a gate-defined QD device that illustrates the critical features of accumulation-mode, undoped devices: multiple channels (three white arrows), regions for forming quantum dots (four in the qubit channel, below, and two single-QD charge sensors, above), and reservoir gates to connect these regions to ohmic contacts for current flow. This device exemplifies a commonly used fabrication strategy known as overlapping gates [27, 28].
@@
The screening gates, shown in yellow, are the closest layer of gates to the top gate oxide in the three-layer gate design and are used to isolate each of the channels from each other and from other regions in the chip. The second layer (shown in blue and red) includes the reservoir and plunger gates. The reservoirs enable the flow of electrons from the ohmic contacts toward the center of the device where QDs will form in the electrostatic potential profile controlled by each plunger gate. The final layer consists of barrier gates (grey), which control the tunnel couplings between QDs—both qubit QDs and charge-sensing QDs—and between the outermost QDs and reservoirs. By carefully adjusting the voltages applied to each of the gates, the quantum dot qubit chip with charge sensors is made functional [2, 3, 4, 5].
@@
BATIS is designed to automate the task of testing and characterizing all gates and forming all necessary 1D current channels for arbitrary-length linear QD arrays. The procedural flow of the single button press algorithm is depicted in Fig. 1(b).
@@
BATIS takes as input two configuration files describing the device design and certain generic characteristics of the device and instruments. Based on that information, it performs full diagnostics and characterization of each gate and determines the proper configuration for the consecutive stage of the tuning processes. Each of the six stages in BATIS involves at least one device measurement followed by data analysis to test the functionality of a selected subset of gates, shown in Fig. 1(b) by gray rounded rectangles. The optional illumination stage is executed on an as-needed basis. Illumination involves exposing the QD device to 780 light under carefully chosen gate voltages, to modify its state. Such exposure shifts the gate voltage response by modifying the amount (and if desired, the sign) of charge trapped at the semiconductor-oxide interface [26].
@@
The algorithm follows a conditional logic visually represented with arrows of different colors and textures in Fig. 1(b). The solid gray arrows indicate the procedural step in the sequence. Passing a test for a current stage, marked with yellow diamonds in Fig. 1(b), initiates a successive measurement and assessment stage, as indicated with solid green arrows. Failing a test can result in repeated measurements, device illumination, or tuning termination, depending on the stage. The measurement schemes, together with idealized and real-world data associated with each stage, are shown in Fig. 2.
@@
Once completed, BATIS produces a device diagnostics file that stores all information about each gate learned during tuning. It also pre-configures the QD device for further use. The details of the flow of the algorithm are discussed in the following sections.
@@ II.1 Initialization and First Leakage Test
The two configuration files used as input for BATIS include descriptions of all gates, desirable current channels, and the QD device connections. The files include also details about the software configuration necessary to execute BATIS and a list of variables that need to be considered at each stage. These variables consist of instrumentation limits to ensure that the algorithm does not search in ranges not supported by the instruments and device limits to prevent the device from breaking during tune-up. Details included in configuration files allow the linking of abstract concepts about the device components with the language used in BATIS.
The functionality of gates is determined automatically from the device architecture that the end user inputs, building a gate-lookup table. The files also contain a wire map and instrument information detailing the measurement system setup such that BATIS can automatically read and execute all specialized measurements, connecting the gate-lookup table with the measurement instruments. This allows BATIS to remain device- and instrument-agnostic.
BATIS starts by checking the device for leakage to the fridge ground through the oxide. The first leakage measurement happens before the electron accumulation during the leakage test stage, see Fig. 1(b). The resistances of each connection are measured with respect to each other and compiled into a leakage matrix, as shown in the second and third row of Fig. 2(a). The desired result is a high resistance between all the connections, an entirely yellow matrix.
@@ II.2 Global Accumulation
Idealized successful and failing measurements for this stage are visualized in the left and right panels in the middle row of Fig. 2(b), respectively. The turn-on process in real QD devices is not always straightforward, as even a good device might not turn on properly when first tested. If the algorithm encounters a problem within global turn-on, it has a recovery pathway through optional illumination, see Fig. 1(b).
@@ II.3 Accumulated Leakage Test
Once the 2DEG is accumulated, the second, accumulated leakage test is initiated. This step proceeds analogously to the first leakage check, except that it does not include the ohmic connections. This is because this leakage check is taken when the device is accumulated, and the ohmics are intentionally shorted to each other. If this stage fails, it indicates that a gate is shorted to the 2DEG flowing beneath it, which is an unrecoverable fabrication defect. Should that occur, BATIS terminates.
@@ II.4 Screening and Reservoir Characterization
The turn-on characteristic voltage is taken globally, meaning it does not provide gate-specific information. To learn that information, BATIS performs a series of individual gate characterizations. The first of these is performed at the screening and reservoir characterization stage in Fig. 1(b).
With the 2DEG accumulated and the device turned on, BATIS depletes the voltages on individual gates until the current on the corresponding channel is pinched off, as displayed in the right panel in the second row in Fig. 2(c). BATIS uses these pinch-off curves to determine where to place reservoir gates for proper 2DEG accumulation and establish bounds to operate the screening gates.
The typical failure modes for this stage include too much depletion and too much accumulation. If the channel below a gate is extraordinarily depleted, no voltage can be applied to the gate to accumulate it. However, since the screening and reservoir characterization stage is only initialized if all current channels turn on properly at the global accumulation stage, the inability to accumulate implies either an issue with the measurement instrument or hysteresis with the gate’s action upon the 2DEG. Similarly, pinching off of all current channels is also confirmed at the global accumulation stage, and thus, it is implausible to observe too much accumulation at this stage. To test whether the issue is caused by the instrument, the algorithm executes a repeated measurement. If the error is observed a second time, it indicates an issue with the gate and the algorithm terminates.
@@ II.5 Forming 1D Current Channels
Selectively removing 2DEG from certain regions of the device is done to form the 1D current channels. It requires depleting the screening gates and accumulating the plungers and barriers (often collectively called finger gates) such that the current flows between the screening gates along a desirable path, see the left panel in Fig. 2(c).
Determining voltages where 1D channels are properly formed is typically done by sweeping two screening gates simultaneously in a 2D sweep while holding the finger gates constant. The formation of the channel manifests as a triangular-sloped region (the so-called triangle plots). If the finger gates are not accumulated enough, this region will not form. Thus, multiple scans with incremented finger gate voltages—a three-dimensional dataset—are often required before the region is observed for all required channels.
BATIS employs a different approach to simultaneously accumulate multiple 1D channels using only a single two-dimensional dataset. Instead of sweeping two screening gates, BATIS takes a 2D voltage scan while sweeping one screening gate against all of the finger gates between that gate and the other screening gate used to form the 1D channel (typically the central screening gate). Since BATIS already learned the necessary configuration for the channel isolation with the central screening gate, this information is now utilized to configure the remaining screening gates. The 1D channel formation technique implemented in BATIS requires a single measurement for each pair of gates, which makes it much more efficient than the standard approaches.
The plot in the right panel in the second row Fig. 2(d) exhibits three characteristic regions: a no-current region in the bottom left, a screening-gate-dominated region in the bottom right, and a finger-gate-dominated region in the top left. BATIS automatically selects an operating point in the top left part of the plot, near the boundary between the current and the no-current regions. The automatic analysis of the regions and their respective boundaries is enabled by fitting a synthetic model to the experimental data, similar to the presentation in [29]. Additional criteria based on the local properties, e.g., estimated resistance and derivatives of the measurement, are used to optimize the selection of the final point among numerous candidates in the operating region.
If the finger-gate-dominated region does not form, BATIS illuminates the device with the finger gates at a voltage 400 more negative than the most recent illumination. After such an illumination, the entire device response can change, because of cross-talk and the interdependence of the current paths. Thus, BATIS resumes autotuning from the global accumulation step.

### v2

@@ Abstract
Semiconductor quantum dot (QD) devices have become central to advancements in spin-based quantum computing. As the complexity of QD devices grows, manual tuning becomes increasingly infeasible, necessitating robust and scalable autotuning solutions. Tuning large arrays of QD qubits depends on efficient choices of automated protocols. Here, we introduce a bootstrapping, autonomous testing, and initialization system (BATIS) designed to streamline QD device evaluation and calibration. BATIS navigates high-dimensional gate voltage spaces, automating essential steps such as leakage testing and gate characterization. For forming the current channels, BATIS follows a non-standard approach that requires a single measurement regardless of the number of channels. Demonstrated at on a quad-QD Si/SixGe1-x device, BATIS eliminates the need for deep cryogenic environments during initial device diagnostics, significantly enhancing scalability and reducing setup times. By requiring only minimal prior knowledge of the device architecture, BATIS represents a platform-agnostic solution, adaptable to various QD systems, which bridges a critical gap in QD autotuning.
@@ I Introduction
Quantum dot (QD) qubit devices are commonly formed by electrostatically manipulating the potential energy landscape of a two-dimensional (2D) electron or hole gas (EG or HG). This procedure produces interleaved regions of accumulated and depleted charge density that serve as QD qubits, charge sensors, and charge carrier reservoirs [1, 2, 3, 4, 5]. In recent years, the number of QDs in such devices has risen rapidly [6, 7, 8]. In the near term, it will no longer be possible to tune up many dots simultaneously using human-based techniques. The field of autotuning has arisen to address this challenge.
@@
Modern semiconductor QD qubit devices have two critical features that enhance their performance, while simultaneously making autotuning challenging and essential. First, the active region of the device—the QDs, charge sensors, and reservoirs—contains no doping. All charge carriers—electrons or holes—are induced by gate voltages. No current can flow until gate voltages are applied and carriers are accumulated, yet turn-on voltages are unknown a priori. Furthermore, because of device drift and hysteresis, the exact voltage configuration changes not only between cooldowns but even between successive initialization steps. Second, QD device chips have multiple current channels that often serve different purposes. Some can host an array of QD qubits, while others provide a home for charge sensors. These channels are not separate until pinch-off voltages, i.e., the gate voltage threshold at which carrier accumulation is eliminated, are determined 111For a given current curve, the pinch-off and the turn-on voltage are mathematically indistinguishable. However, when measuring a physical gate, a hysteresis (i.e., a shift of the current curve that depends on the direction of the sweep) can be observed.. Moreover, since the accumulation and depletion of carriers must occur in alternating regions separated by only tens of nanometers, these channels must interact to enable the formation of QDs separated by tunnel barriers and proximal charge sensors.
@@
Here, we describe a bootstrapping, autonomous testing, and initialization system (BATIS) that can automatically manipulate gate voltages to ensure that the QD chip is properly accumulated and all required pinch-off voltages are determined. BATIS takes as input the device gate layout and a high-level understanding of the desired operation, including (critically) where current channels should be continuous and which gates should divide one channel from another. The outputs of BATIS are the gate voltages that make the QD chip functional, with all current channels well delineated and all pinch-off voltages known. Importantly, in this work, the sample is kept at around —about 10 times warmer than other typical devices [10, 11]—which is highly desirable for high-temperature operation [12]. However, higher temperatures broaden key measurement features, such as Coulomb blockade peaks, necessitating careful fitting procedures to extract the desired parameters [13]. We show that BATIS works well at these high temperatures, which is important both for scalability and automated testing, e.g., using a cryogenic wafer prober [14, 15].
@@
Previous work on autotuning has shown that correctly executed bootstrapping of QD devices sets the stage for the more specialized tuning phases. For example, setting the device topology in a modulation-doped QD device with predefined current channels has been demonstrated in Refs. [16, 17, 18]. Autotuning has similarly been implemented in an undoped SiMOS QD to achieve single-electron occupation [19]. For nanowires, which by design have a single current channel, automated tuning into fully formed QDs has been demonstrated, provided device-specific details were preconfigured [20].
@@
Additionally, for undoped, 1D accumulation-mode QD devices, previous work has enabled progress in tuning a device topology [22, 23] and calibrating the charge state [16, 24]. Methods for automatically detecting unintentional, spurious QDs have also been proposed [25]. More recently, virtual-gate control of the electrostatic potential landscape confining holes within a 2D QD array has been demonstrated [26]. However, autonomous bootstrapping for accumulation-mode QD devices has never been attempted. Rather, most proposed tuning algorithms assumed that initial steps, such as device characterization and channel formation, have already been accomplished [27].
Our approach for BATIS differs from the existing bootstrapping algorithms, which often use efficiency as a figure of merit. By focusing on speed, such algorithms tend to be specialized to a particular device, allowing little room for benchmarking on different devices or in different experimental setups. Even if such algorithms perform well in one setting, they require a significant redesign to work in other settings. In contrast, we have specifically designed BATIS to be platform-agnostic, so that it may be applied to a variety of QD devices. While the procedural flow of BATIS is well-defined, the executable flow is automatically adjusted to suit the device architecture and external control lines, both of which are specified in the configuration file. In other words, rather than following a fixed sequence of measurements, BATIS—in real-time—chooses which steps to take and defines the expected outcomes depending on the device used. As the tuning progresses and the algorithm learns the device’s characteristics, BATIS applies its knowledge to automatically customize further measurements.
Finally, we note that existing autotuning schemes focus mainly on the more advanced stages of a tune-up. These include coarse tuning, charge-state tuning, virtualization, and fine-tuning, which are sufficient for QDs formed in a preexisting channel, or when the initial bootstrapping steps are performed by hand [27]. In modern chips, hand-tuning will soon be precluded as QD processors scale up. BATIS is designed to autonomously adjust to the QD layout and perform the initial tune-up stages, including checking for fabrication defects and characterizing local and global current flow, all of which are essential for large-scale qubit implementations. It thus addresses the challenges that are about to become limiting in this widespread class of quantum semiconductor chips.
@@ II BATIS Design
Figure 1(a) shows a gate-defined QD device that illustrates the critical features of accumulation-mode, undoped devices: multiple channels (three white arrows), regions for forming QDs (four in the qubit channel, below, and two single-QD charge sensors, above), and reservoir gates to connect these regions to ohmic contacts for current flow. This device exemplifies a commonly used fabrication strategy known as overlapping gates [28, 29].
@@
The screening gates (yellow) are the closest layer of gates to the top gate oxide in the three-layer gate design and screen electric fields to prevent unwanted regions of 2DEG/2DHG, effectively isolating the channels. The second layer includes the reservoir (blue) and plunger (red) gates. The reservoirs enable the flow of electrons from the ohmic contacts toward the center of the device, where QDs will form in the electrostatic confined potential. The chemical potential of each QD is primarily controlled by the corresponding plunger gate. The final layer consists of barrier gates (grey), which control the tunnel couplings between QDs—both qubit QDs and charge-sensing QDs—and reservoirs. By carefully adjusting the voltages applied to each of the gates, the QD qubit chip with charge sensors is made functional [2, 3, 4, 5].
@@
BATIS is designed to automate the task of testing and characterizing all gates and forming all necessary 1D current channels for arbitrary-length linear QD arrays. The procedural flow of the single-button-press algorithm is depicted in Fig. 1(b).
@@
BATIS takes as input two configuration files describing the device design and certain generic characteristics of the device and instruments. Based on that information, it performs full diagnostics and characterization of each gate and determines the proper configuration for the consecutive stage of the tuning processes. Each of the six stages in BATIS involves at least one device measurement followed by data analysis to extract characteristics of a selected subset of gates, shown in Fig. 1(b) by gray rounded rectangles. The optional optical illumination stage is executed on an as-needed basis. Illumination involves exposing the QD device to 780 light under carefully chosen gate voltages, to modify its state. Such exposure shifts the gate voltage response by modifying the amount (and if desired, the sign) of charge trapped at the semiconductor-oxide interface [21].
@@
The algorithm follows conditional logic visually represented with arrows of different colors in Fig. 1(b). The gray arrows indicate the procedural step in the sequence. Passing the test for the current stage, marked with yellow diamonds in Fig. 1(b), initiates the next stage, as indicated with green arrows. Failing a test can result in repeated measurements, device illumination, or termination of tuning, depending on the stage. The physical intuition behind each measurement, together with idealized and real-world data associated with each stage, are shown in Fig. 2.
@@
Once completed, BATIS produces a device diagnostics file that stores all information about each gate learned during tuning. It also preconfigures the QD device for further use. The details of the flow of the algorithm are discussed in the following sections.
@@ II.1 Initialization and First Leakage Test
The two configuration files used as input for BATIS include descriptions of all gates, desirable current channels, and the QD device connections. The files also include details about the software configuration necessary to execute BATIS and a list of variables that need to be considered at each stage. These variables consist of instrumentation limits to ensure that the algorithm does not search in ranges not supported by the instruments and device limits to protect the device from having large damaging voltages applied. Details included in configuration files allow the linking of abstract concepts about the device components with the language used in BATIS.
For example, the functionality of gates is determined automatically from the device architecture that the end user inputs, building a gate-lookup table. The files also contain a wire map and instrument information detailing the measurement system setup such that BATIS can automatically read and execute all specialized measurements, connecting the gate-lookup table with the measurement instruments. This allows BATIS to remain device- and instrument-agnostic.
BATIS starts by checking the device for leakage to the fridge ground through the oxide or any possible means in the instrumentation setup. The first leakage measurement happens before the electron accumulation during the leakage test stage, see Fig. 1(b). The resistances of each connection are measured with respect to each other and compiled into a leakage matrix, as shown in the second and third row of Fig. 2(a). The desired result is a high resistance between all the connections.
@@ II.2 Global Accumulation
Determining the turn-on voltages VP relies on fitting to a phenomenological function that mimics the expected behavior of the signal. Prior to fitting, the raw data is preprocessed to clean the signal. The preprocessing involves cropping the data to avoid non-linearities at the extreme limits of the data acquisition electronics followed by smoothing the data to remove noise from the signal. The processed signal is then fit to a function of the form
where is the center of the signal and describes the slope at . The final turn-on voltage is defined as
where is a fixed constant specifying how much the gate needs to be adjusted to turn-on in the units of 222The parameter is defined as part of the configuration file and can be adjusted if needed.. In addition, the saturation of the curve is defined following Eq. (2), with . In cases where the saturation of the curve is not captured in the measurement, the last point along the sweep axis becomes . The first estimation of the maximum allowed voltage applied to each gate is defined as V.
Idealized successful and failing measurements for this stage are visualized in the left and right panels in the middle row of Fig. 2(b), respectively. The turn-on process in real QD devices is not always straightforward, as even a good QD device might not turn on properly when first tested. If the algorithm encounters a problem within global turn-on, it has a recovery pathway through optional illumination, see Fig. 1(b).
@@ II.3 Accumulated Leakage Test
Once the 2DEG is accumulated, the accumulated leakage test is initiated. This step proceeds analogously to the first leakage check, except that it does not include the ohmic connections. This is because this leakage check is taken when the device is accumulated, and the ohmics are intentionally shorted to each other. If this stage fails, it indicates that a gate is shorted to the 2DEG flowing beneath it, which is an unrecoverable fabrication defect. Should that occur, BATIS terminates.
@@ II.4 Screening and Reservoir Characterization
The global accumulation voltage does not provide gate-specific information. To learn that information, BATIS performs a series of individual gate characterizations. The first of these is performed at the screening and reservoir characterization stage in Fig. 1(b).
With the 2DEG globally accumulated using the voltages selected from Sec. II.2, BATIS depletes the voltages on individual gates until the current on the corresponding channel is pinched off, as displayed in the left panel in the second row in Fig. 2(c). BATIS uses these pinch-off curves to determine the operational configuration for reservoir gates that ensures proper 2DEG accumulation and establishes bounds for operating the screening gates.
The signals for screening and reservoir gates are preprocessed as described in Sec. II.2. Once cleaned, the derivative of the current vs. voltage signal, , is fit to the derivative of the phenomenological function given in Eq. (1). The pinch-off, guaranteed isolation, and operating voltages are defined following Eq. (2), with , , and , respectively. The operating voltage for the central screening gate is handled separately as it has to satisfy two constraints: ensuring that all current channels can be formed without merging. By default, . If at the subsequent state, a failure to form 1D current channels is observed twice, BATIS adjusts this parameter in increments defined as , where is chosen at random.
The typical failure edge cases for this stage include too much depletion and too much accumulation; see the second row, the right column of Fig. 2(c). If the channel below a gate is extraordinarily depleted, no voltage can be applied to the gate to accumulate it. On the other hand, if the channel is too accumulated, it is not possible to pinch it off. However, both types of issues would have been detected at the global accumulation stage, prompting the initiation of device illumination. Thus, the no change in current at this stage has to be caused either by an issue with the instrument or a broken gate. To check whether this is the case, the algorithm executes a repeated measurement. If the error is observed a second time, it indicates an issue with the gate and the algorithm terminates.
@@ II.5 Forming 1D Current Channels
The 1D current channels are formed by selectively removing 2DEG from certain regions of the device. It requires depleting the screening gates and accumulating the plungers and barriers (often collectively called finger gates) such that the current flows between the screening gates along the desirable path, as seen in the top row of Fig. 2(d).
Determining voltages where 1D channels are properly formed is typically done by sweeping two screening gates simultaneously in a 2D sweep while holding the finger gates constant. The formation of the channel manifests as a triangular-sloped region (the so-called triangle plots) [31]. If the finger gates are not accumulated enough, this region will not form. Thus, multiple scans with incremented finger gate voltages—a three-dimensional dataset—are often required before the region is observed for all required channels.
BATIS employs a different approach to simultaneously accumulate multiple 1D channels using only a single two-dimensional dataset. Instead of sweeping two screening gates, BATIS takes a 2D voltage scan while sweeping one screening gate against all of the finger gates between that gate and the other screening gate used to form the 1D channel (typically the central screening gate). BATIS uses the channel isolation (VI) and the operating points (VO) to prepare the device for characterizing the remaining screening gates by sweeping from VI past VP. The 1D channel formation technique implemented in BATIS requires a single measurement for each pair of gates, which makes it much more efficient than the standard approaches.
The plot in the left panel in the second row Fig. 2(d) exhibits three characteristic regions: a no-current region in the bottom left, a screening-gate-dominated region in the bottom right, and a finger-gate-dominated region in the top left. BATIS automatically selects an operating point in the top left part of the plot, near the boundary between the current and the no-current regions. The automatic analysis of the regions and their respective boundaries is enabled by fitting a synthetic model to the experimental data, similar to Ref. [32]. Additional criteria based on the local properties, e.g., estimated resistance and derivatives of the measurement, are used to optimize the selection of the final point V for , among numerous candidates in the operating region.
If the finger-gate-dominated region does not form, as shown in the right panel in the second row of Fig. 2(d), BATIS illuminates the device with the finger gates voltages set to be more negative than the most recent illumination. This adjustment can be found in the configuration file, we used 400 . After such an illumination, the entire device response can change, because of cross-talk and the interdependence of the current paths. Thus, BATIS resumes autotuning from the global accumulation step.